OpenStudy (anonymous):

If the arc on a particular circle has an arc length of 3 feet, and the angle measure of the arc is 60°, what is the circumference of the circle?

OpenStudy (mertsj):

Would you agree that if the arc is 60 degrees that is 1/6 of the entire circle? So perhaps, is you know 1/6 of the circle, you can find 6/6 of the circle.

OpenStudy (gorv):

use l=r*theta l=arc length r=radius 3=r*pi/3 r=9/pi circumference=2pi*r =2*pi*9/pi =18